# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v1

@@ 4.1 Estimating $r^{2}$, $\sigma^{2}$, and $\alpha_{i}$
4:$\hat{\boldsymbol{\alpha}}\leftarrow(\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{X}_{\mathsf{proj}})^{-1}\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{y}$
@@
7:return $\mathcal{R}_{\mathbf{X},\mathbf{y}}\left(\hat{r}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}}),\hat{\sigma}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}}),\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}}\right)$
Our approach is simply to estimate the $\alpha_{i}$’s using principle components regression (see [25, 23, 4, 22, 10, 2, 21, 47, 24, 35, 37] for details on principle components regression), described in steps 1 and 2 below. We then compute a reduced model which is right-rotationally invariant, from which vanilla $\mathsf{ROTI}\text{-}\mathsf{GCV}$ can be used to find $r^{2}$ and $\sigma^{2}$.
Throughout all that follows, we focus on subsets $S$ of the indices of all nonzero singular values $\mathcal{N}=\{i\in[p]:(\mathbf{D}^{\top}\mathbf{D})_{ii}>0\}$. Let $\mathbf{O}_{S}\in\mathbb{R}^{|S|\times p}$ for any set $S$ denote the rows of $\mathbf{O}$ indexed by $S$. Let $\mathbf{D}_{S}\in\mathbb{R}^{(n-|S|)\times p}$ denote the rows of $\mathbf{D}$ indexed by $S$. Finally, for a set $S\subset\mathcal{N}$, denote $\overline{S}=\mathcal{N}\setminus S$. The procedure then proceeds as in steps 1-4 of Algorithm 2.
Lemma 5.
Under the additional assumption that $\boldsymbol{\epsilon}$ has Gaussian entries, as described above, $\hat{r}^{2}$, $\hat{\sigma}^{2}$, and $\hat{\boldsymbol{\alpha}}$ are all strongly consistent.
The assumption of Gaussian entries is technical and we believe it can be removed with some work.
4.2 Cross-validation metric
Our approach to forming a cross-validation metric is to then again explicitly compute the risk formulas
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)}{(D_{ii}^{2}+\lambda)^{2}},$
$\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\sigma^{2}}{n}\left[\sum_{i=1}^{n}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)\right],$
where we estimate the unknown parameters from data. To apply our method, we therefore need to consistently estimate $r^{2},\sigma^{2},\boldsymbol{\alpha}$, and $\{\mathfrak{d}_{i}^{2}\}_{i\in\mathcal{J}_{c}}$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}$. Lemma 5 gives us most of these parameters, but we still require consistent estimates of the $\mathfrak{d}_{i}$’s. This is a nonissue in cases where $\mathbf{D}$ and $\widetilde{\mathbf{D}}$ are modeled as deterministic. If they are random, we expect that if the test data is drawn from the same right-rotationally invariant ensemble as the training data, then, in practice, one can use the eigenvalues of the training data to do this (for theoretical guarantees under some examples of random $\mathbf{D}$, see results on eigenvalue rigidity [9]). However, our framework further allows for the test distribution to be drawn from a different right-rotationally invariant ensemble. In such cases, it may be possible to then use test data to do this estimation because it only depends on the test covariates $\widetilde{\mathbf{X}}$ (which the practitioner may have access to), but not its labels. We therefore make one final assumption:
  1. A11
We then take our validation metric as $\mathsf{aROTI}\text{-}\mathsf{GCV}(\lambda)=\mathcal{R}_{\mathbf{X},\mathbf{y}}(\hat{r}^{2},\hat{\sigma}^{2},\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}_{i}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}})$, where the latter expression denotes substituting every parameter with its estimated version.
@@ 1. A11
Under Assumptions A1-A4, A7-A11, and the additional assumption that the entries of $\boldsymbol{\epsilon}$ are Gaussian, $\mathsf{aROTI}\text{-}\mathsf{GCV}$ satisfies
@@
The total algorithm was described above in Algorithm 2.
@@
[27] presents a way to do hypothesis testing for alignment to detect $\mathcal{J}_{a}$. For detecting $\mathcal{J}_{c}$, one can measure the angle among the top eigenvectors of the test set and top eigenvectors of the training set, placing pairs of vectors which have overlap deviating heavily from $90^{\circ}$ into $\mathcal{J}_{c}$.
@@ 5 Numerical Experiments
5.1 Performance of ROTI-GCV on right-rotationally invariant data
Figure 1 illustrates the performance of ROTI-GCV and compares it to LOOCV and GCV. In these figures, tuned risk refers to the out-of-sample prediction error obtained when we use the value of $\lambda$ that optimizes the given cross-validation method. For the Expected MSE, we report the minimum of the true expected out-of-sample mean squared error curve. Standard errors are shown for all tuned risks. Est Risk refers to the value of the out-of-sample risk that would be estimated using the given cross-validation method. As expected, in the Gaussian setting all three methods perform well. However, for dependent data, such as in the autocorrelated and equicorrelated cases, we observe that the risk curves produced by GCV and LOOCV are wildly off from the truth, whereas ROTI-GCV accurately captures the true risk curve. Surprisingly, the tuned risk values from LOOCV, GCV are still decently close to that of ROTI-GCV. Because these two methods estimate the total mean squared error $\mathbb{E}[(y-\mathbf{x}^{\top}\hat{\boldsymbol{\beta}})^{2}\mid\mathbf{X},\mathbf{y}]=R_{\mathbf{X}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})+\sigma^{2}$ , the values they produce include an additional factor of the noise level $\sigma^{2}$, while the risk we are interested (See equation (2)) does not include this term, and hence are plotted with this term removed.
@@ 5.1 Performance of ROTI-GCV on right-rotationally invariant data
Figure 1: Risk curves produced by the cross-validation methods (LOOCV, GCV, ROTI-GCV), in addition to the true risk curve. Each plot uses a different form of right-rotationally invariant data, as indicated in each caption. (b): Autocorrelated data: rows of $\mathbf{X}$ are drawn according to $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\boldsymbol{\epsilon}_{i}$, with $\boldsymbol{\epsilon}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$. We set $\rho=0.8$. (c): $t$-distributed data: each row is drawn from a multivariate $t$ distribution with $3$ degrees of freedom. (d): Equicorrelated data $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$. All simulations have $n=p=1000$ and $r^{2}=\sigma^{2}=1$. The $x$-axis for every plot is $\lambda$, the regularization parameter. The colored lines (blue, green, red, purple) are one of 10 iterations. In each iteration, we compute the cross-validation metric over a range of $\lambda$ to produce the line, which reflects the estimated out-of-sample risk. Theoretical risk is plotted in all three as reference (black line), and refers to the exact value of $R_{\mathbf{X}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})$ (see equation 2). The dashed blue line shows the average expected MSE curve as a benchmark.
@@ 5.2 Experiments under Signal-PC alignment
As discussed in Section 4, the scenario where the signal $\boldsymbol{\beta}$ aligns with the eigenvectors $\mathbf{O}$ is problematic for right-rotationally invariant designs. In Figure 2, we show that the aROTI-GCV method from Section 4 works when our assumptions are met in a somewhat contrived example where we set the top eigenvectors of the test set to be identical to those of the training set, matching our coupled eigenvectors condition. In Figure 3, we give some examples in which this situation can arise.
Figure 2 illustrates the accuracy of our method when this alignment is introduced. Here, the signal aligns with the top 10 eigenvectors, and these top 10 eigenvectors are also coupled with the test set. As expected, the theory matches when these assumptions are met.
Figure 2: GCV under Signal-PC alignment. As in Figure 1, curves show the risk and estimated risk as a function of $\lambda$. $\mathbf{X}$ is drawn with autocorrelated rows. We set $\rho=0.8$. The top 10 eigenvectors ($\mathcal{J}_{c}=[10]$) are coupled, so that the top 10 eigenvectors of the test set are equal to those of $\mathbf{X}$. $\boldsymbol{\beta}$ is drawn according to $\boldsymbol{\beta}^{\prime}+\sqrt{n}\sum_{i=1}^{10}\frac{i}{10}\mathbf{o}_{i}$, with $\|\boldsymbol{\beta}^{\prime}\|^{2}=1$, $\sigma^{2}=1$, $n=p=1000$.
@@
(a) i.i.d. rows drawn from a Gaussian mixture, i.e. $\mathbf{x}_{i}$ is drawn from $\frac{1}{2}\mathsf{N}(\mathbf{3},\mathbf{I}_{p})+\frac{1}{2}\mathsf{N}(-\mathbf{3},\mathbf{I}_{p})$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=\{1\}$.
(b) i.i.d. rows drawn from an equicorrelated Gaussian, i.e. $\mathbf{x}_{i}$ is drawn from $\mathsf{N}(0,\boldsymbol{\Sigma})$, where $\Sigma_{ij}=\mathds{1}(i=j)+\rho\mathds{1}(i\neq j)$. We set $\rho=0.5$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=[10]$.
Figure 3: GCV performance on data with correlation. As in Figure 1, curves show the risk and estimated risk as a function of $\lambda$.
@@ 6 Discussion and Conclusion
Takeaways: We introduce ROTI-GCV, a method for estimating out-of-sample prediction error under a new model for data, namely right-rotationally invariant covariates. The model departs crucially from the existing framework of having i.i.d. samples drawn from a given distribution and sub-Gaussian tailed covariates. Instead it allows for structured forms of dependence and heavy tails. ROTI-GCV is a new cross-validation framework that is provably consistent for these designs and we demonstrate its superiority over existing methods across diverse settings. As a byproduct of our analysis, we produce new estimators for signal-to-noise ratio and noise variance that are robust to these deviations from the i.i.d. assumptions. At the core, our idea involves identifying the relation given by Lemma 2 and noticing that the term on the right involves $\lambda$. This means we can iterate and compute this term over a grid of values of $\lambda$. This leads to a series of estimating equations from which stable estimators of $r^{2},\sigma^{2}$ can be derived. This core idea can be applicable for choosing hyperparameters for much broader models beyond linear regression–pretty much anywhere an analogue of Lemma 2 is available. As future steps, extending our method to other settings such as generalized linear models, kernel ridge regression, regularized random features regression would be important directions. Furthermore, recent advances in [8, 44] show that right-rotationally invariant distributions have a broad universality class–in fact as long as the singular vectors of a design matrix satisfy certain generic structure, properties of learning algorithms under such settings are well captured by results derived under the right-rotationally invariant assumption. In particular, [8, 44] establish such universality properties of the in-sample risk for ridge regression. We expect analogous universality results to hold in our setting. Thus, ROTI-GCV should apply broadly beyond right-rotationally invariant designs. In fact, it should be effective for any design that belongs to the spectral universality class identified by [8]. A comprehensive exploration of how this universality can be used to improve hyperparameter tuning is a promising direction for future research.
Limitations: Right-rotationally invariant designs may on occasion have difficulty modeling distributions with feature covariances that are much more complicated than those covered in Section 5, especially when the spectrum is not dominated by a few top eigenvectors. Finding ways to account for different types of feature covariance in data using these designs would naturally increase the broader applicability of our method.
7 Acknowledgements
P.S. would like to acknowledge partial funding from NSF DMS-2113426.

### v2

@@ 4.1 Estimating $r^{2}$, $\sigma^{2}$, and $\alpha_{i}$
Our approach is simply to estimate the $\alpha_{i}$’s using the classical principal components regression (PCR);
see [20, 19] for details, also described in steps 1 and 2 of Algorithm 2. We then compute a transformed model which is right-rotationally invariant, from which Algorithm 1 steps 1-6 can be used to find $r^{2}$ and $\sigma^{2}$.
We require some notation; we focus on subsets $S$ of the indices of all nonzero singular values $\mathcal{N}=\{i\in[p]:(\mathbf{D}^{\top}\mathbf{D})_{ii}>0\}$. Let $\mathbf{O}_{S}\in\mathbb{R}^{|S|\times p}$ for any $S$ denote the rows of $\mathbf{O}$ indexed by $S$. Let $\mathbf{D}_{S}\in\mathbb{R}^{(n-|S|)\times p}$ denote the rows of $\mathbf{D}$ indexed by $S$. Finally, for a set $S\subset\mathcal{N}$, denote $\overline{S}=\mathcal{N}\setminus S$. The estimation for $r^{2},\sigma^{2}$, and $\boldsymbol{\alpha}$ then proceeds as in steps 1-5 of Algorithm 2.
Lemma 4.
Under the additional assumption that $\boldsymbol{\epsilon}$ has Gaussian entries, $\tilde{r}^{2}$, $\tilde{\sigma}^{2}$, and $\hat{\boldsymbol{\alpha}}$, from Algorithm 2, are all strongly consistent.
The assumption of Gaussian entries is technical, and we believe it can be removed with some work.
4.2 Cross-validation metric
Our approach for cross-validation is to then again explicitly compute the risk formulas $\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ and $\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$, where we estimate the unknown parameters from data. To apply our method, we therefore need to consistently estimate $r^{2},\sigma^{2},\boldsymbol{\alpha}$, and $\{\mathfrak{d}_{i}^{2}\}_{i\in\mathcal{J}_{c}}$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}$. Lemma 4 gives us most of these parameters, but we still require consistent estimates of the $\mathfrak{d}_{i}$’s. This is a nonissue in cases where $\mathbf{D}$ and $\widetilde{\mathbf{D}}$ are modeled as deterministic. If they are random, we expect that if the test data is drawn from the same right-rotationally invariant ensemble as the training data, then, in practice, one can use the eigenvalues of the training data to do this. However, our framework further allows for the test distribution to be drawn from a different right-rotationally invariant ensemble. In cases where test data covariates are available (we do not require access to the labels), it is possible to then use test data to do this estimation; this is done for experiments in Section 5. We therefore make one final assumption:
  1. A11
We then take our validation metric as $\mathsf{aROTI}\text{-}\mathsf{GCV}(\lambda)=\mathcal{R}_{\mathbf{X},\mathbf{y}}(\tilde{r}^{2},\tilde{\sigma}^{2},\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}_{i}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}})$, where the latter expression denotes substituting every parameter with its estimated version. Algorithm 2 summarizes our entire procedure.
@@ 1. A11
4:$\hat{\boldsymbol{\alpha}}\leftarrow(\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{X}_{\mathsf{proj}})^{-1}\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{y}/\sqrt{n}$
@@
7:Compute $\tilde{r}=\hat{r}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}}),\tilde{\sigma}^{2}=\hat{\sigma}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}})$
8:return $\mathcal{R}_{\mathbf{X},\mathbf{y}}\left(\tilde{r}^{2},\tilde{\sigma}^{2},\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}}\right)$
@@
Under Assumptions A1-A5, A8-A11, and the additional assumption that the entries of $\boldsymbol{\epsilon}$ are Gaussian, $\mathsf{aROTI}\text{-}\mathsf{GCV}$ satisfies
@@
Remark 7.
This states that aROTI-GCV is uniformly consistent on compact intervals, again ensuring that the minimizer of aROTI-GCV attains risk that is close to optimal.
@@
[23] presents a way to do hypothesis testing for alignment to detect $\mathcal{J}_{a}$. For detecting $\mathcal{J}_{c}$, one can measure the angle among the top eigenvectors of the test set and top eigenvectors of the training set, placing pairs of vectors which have overlap far from $90^{\circ}$ into $\mathcal{J}_{c}$. We leave the development of a rigorous hypothesis test to future works.
@@ 5 Numerical Experiments
5.1 Experiments on right-rotationally invariant data
@@
Figure 1: Risk curves produced by the cross-validation methods in addition to the true risk curve. Each plot uses a different form of right-rotationally invariant data, as indicated in each caption. (b): Autocorrelated data: rows of $\mathbf{X}$ are drawn according to $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\mathbf{z}_{i}$, with $\mathbf{z}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$. We set $\rho=0.8$. (c): $t$-distributed data: each row is drawn from a multivariate $t$ distribution with $3$ degrees of freedom. (d): Equicorrelated data $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$. All simulations have $n=p=1000$ and $r^{2}=\sigma^{2}=1$. The $x$-axis for every plot is $\lambda$, the regularization parameter. The colored lines (blue, green, red, purple) are one of 10 iterations. In each iteration, we compute the cross-validation metric over a range of $\lambda$ to produce the line, which reflects the estimated out-of-sample risk. The black line of each plot shows the average result for that method. The dashed blue line shows the average expected MSE curve as a benchmark.
Figure 1 illustrates the performance of ROTI-GCV and compares it to LOOCV and GCV. In these figures, the Tuned Risk (abbreviated TR) refers to the out-of-sample risk obtained when we use the value of $\lambda$ that optimizes the given cross-validation method. Minimum risk (MR), denotes the minimum of the true expected out-of-sample risk curve. Standard errors are shown in parentheses for all tuned risks. Estimated Risk (ER) refers to the value of the out-of-sample risk that would be estimated using the given cross-validation method. As expected, in the Gaussian setting all three methods perform well. However, for dependent data, such as in the autocorrelated and equicorrelated cases, we observe that the risk curves produced by GCV and LOOCV are wildly off from the truth, whereas ROTI-GCV accurately captures the true risk curve. Surprisingly, the tuned risk values from LOOCV, GCV are still decently close to that of ROTI-GCV.
As discussed in Remark 3, GCV (and also LOOCV) estimate a quantity with an additional factor of $\sigma^{2}$; thus, to make the evaluation of estimated risks fair, these two curves are plotted with this term removed.
@@ 5.2 Experiments under Signal-PC alignment
(a) Signal-PC alignment
(b) Gaussian mixture
(c) Equicorrelated rows, $\rho=0.5$
(d) Residualized returns
Figure 2: As in Figure 1, curves show the risk and estimated risk as a function of $\lambda$. (2(a)): $\mathbf{X}$ has autocorrelated rows, with $\rho=0.8$. The top 10 eigenvectors ($\mathcal{J}_{c}=[10]$) are coupled, so that the top 10 eigenvectors of the test set are equal to those of $\mathbf{X}$. $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sqrt{n}\sum_{i=1}^{10}\frac{i}{10}\mathbf{o}_{i}$, with $\|\boldsymbol{\beta}^{\prime}\|^{2}=\sqrt{n}$, $\sigma^{2}=1$, $n=p=1000$. (2(b)) i.i.d. rows drawn from a Gaussian mixture, i.e. $\mathbf{x}_{i}$ is drawn from $\frac{1}{2}\mathsf{N}(\mathbf{3},\mathbf{I}_{p})+\frac{1}{2}\mathsf{N}(-\mathbf{3},\mathbf{I}_{p})$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=\{1\}$. (2(c)): i.i.d. rows drawn from an equicorrelated Gaussian, i.e. $\mathbf{x}_{i}\sim\mathsf{N}(0,\boldsymbol{\Sigma})$, where $\Sigma_{ij}=\mathds{1}(i=j)+\rho\mathds{1}(i\neq j)$. We set $\rho=0.5$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=[10]$. (2(d)): 30 minute residualized returns sampled every 1 minute;
We now consider settings where the signal $\boldsymbol{\beta}$ aligns with the eigenvectors $\mathbf{O}$. The relevant method here is aROTI-GCV (Algorithm 2). In Figure 2(a), we show the performance of aROTI-GCV when we set the top eigenvectors of the test set equal to those of the training set, matching the coupled eigenvectors condition (Assumption A9). We further construct the signal such that it aligns with the top 10 eigenvectors. As the conditions match our theory, we see aROTI-GCV performs well. The setting of Figure 2(a) might look contrived, but we next provide examples of well-known distributions where such structures naturally arise (Figures 2(b) and 2(c)).
@@
5.3 Experiments on semi-real data
In our semi-real experiments, we use real data for the designs, but generate the signals and responses ourselves. Knowing the true signal enables us to calculate the population risk of any estimator, allowing us to evaluate the performance of our GCV metrics accurately. The designs in our experiment consist of residualized returns from 493 stocks retrieved from Polygon API [^2] We remove the top 8 right singular vectors from data matrix $\mathbf{X}$. This is analogous to residualization of returns, a common practice in financial modeling. We compute the factors to delete by taking the top PCs of the covariance matrix computed using both test and train data; PCA is crude/weak as a factor model and we have only 8 factors; hence it is more comparable to compute PCs looking at both train and test data. See appendix B.2 for more details on residualization.. There is one row for each minute, and each row contains the residualized return for that stock over the last 30 minutes. As a result of sharing 29 minutes of returns, consecutive rows are heavily correlated. The regression corresponds to the known practice of replicating the residualized returns of an unknown asset $\mathbf{y}$ using a portfolio of stocks $\hat{\boldsymbol{\beta}}$ (with returns $\mathbf{X}\hat{\boldsymbol{\beta}}$). The loss curves produced by ROTI-GCV better capture the true loss curve, and that the estimated risk values (ER) provide superior estimates of the true out-of-sample risk compared to LOOCV and GCV.
@@ 6 Discussion and Conclusion
Takeaways: We give a precise analysis of ridge regression for right-rotationally invariant designs. These designs depart crucially from the existing framework of i.i.d. samples from distributions with light enough tails, instead allowing for structured dependence and heavy tails. We then characterize GCV in this setting, finding that it is inconsistent, and introduce ROTI-GCV, a consistent alternative; various synthetic and semi-real experiments illustrate the accuracy of our method.
As a byproduct of our analysis, we produce new estimators for signal-to-noise ratio ($r^{2}$) and noise variance ($\sigma^{2}$) robust to these deviations from the i.i.d. assumptions. At the core, our method relies on noticing that the relation given by Lemma 1 yields a different estimating equation for every choice of $\lambda$; using this to generate a series of estimating equations, we obtain stable estimators of $r^{2},\sigma^{2}$. This core idea applies whenever an analogue of Lemma 1 can be derived, giving it reach beyond ridge regression. As future steps, extending our method to other settings such as generalized linear models, kernel ridge regression, regularized random features regression would be important directions. Furthermore, recent advances in [8, 41] show that right-rotationally invariant distributions have a broad universality class – in fact as long as the singular vectors of a design matrix satisfy certain generic structure, properties of learning algorithms under such settings are well captured by results derived under the right-rotationally invariant assumption. In particular, the universality of the in-sample risk of ridge regression is already established in these works, and we believe such universality also holds for the out-of-sample risk; thus, our analysis should extend to the entire universality class of these designs, which is remarkably broad. This promises that our approach can have wide reach, allowing reliable hyperparameter tuning for a broad class of covariate distributions.
Limitations: The well-specification assumption (A1) is crucial and seems more difficult to relax for our designs than in the i.i.d. case due to the global dependence structure. Right-rotationally invariant designs have difficulty modeling distributions with feature covariances that are much more complicated than those covered in Section 5, especially when the spectrum is not dominated by eigenvalues. Finding ways to account for different types of feature covariance in data using these designs would naturally increase the broader applicability of our method, such as by modeling the design as $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\boldsymbol{\Sigma}^{1/2}$ and extending our analysis. Finally, as mentioned in Section 1, our results focus on ridge regression, though we believe similar results for other penalties can be obtained.